OpenStudy (anonymous):

Show that if a_n >0 and summation of a_n is convergent, then summation of ln(1+a_n) is convergent

OpenStudy (nowhereman):

Use the majorizing criteria.

OpenStudy (dumbcow):

If sum of a_n converges, then lim a_n =0.(as n goes to infinity) If lim a_n =0, then lim ln(1+a_n)=0. If lim ln(1+a_n)=0, then sum of ln(1+a_n) converges.

OpenStudy (nowhereman):

That is not true. Just knowing that the elements of the series converge to 0 does not mean that the series converges. Take 1/n for example where the series diverges.

OpenStudy (dumbcow):

ok thanks i wasn't sure if my reasoning was correct there. What would be the correct reasoning for showing this? you could show that ln(1+a_n) <= a_n for all n>=0

OpenStudy (nowhereman):

yes, that is true.

OpenStudy (dumbcow):

thanks

OpenStudy (anonymous):

sevenlash, did you get this problem sorted? If not, let me know. I looked at it and have a solution.

OpenStudy (anonymous):

Just not much point writing it out unless you need it :)